# FIKIT: Priority-Based Real-time GPU Multi-tasking Scheduling with Kernel Identification

### 18 Jan 2024 v4

@@ Abstract.
Highly parallelized workloads like machine learning training, inferences and general HPC tasks are greatly accelerated using GPU devices. In a cloud computing cluster, serving a GPU’s computation power through multi-tasks sharing is highly demanded since there are always more task requests than the number of GPU available. Existing GPU sharing solutions focus on reducing task-level waiting time or task-level switching costs when multiple jobs competing for a single GPU. Non-stopped computation requests come with different priorities, having non-symmetric impact on QoS for sharing a GPU device. Existing work missed the kernel-level optimization opportunity brought by this setting. To address this problem, we present a novel kernel-level scheduling strategy called FIKIT: Filling Inter-kernel Idle Time. FIKIT incorporates task-level priority information, fine-grained kernel identification, and kernel measurement, allowing low priorities task’s execution during high priority task’s inter-kernel idle time. Thereby, filling the GPU’s device runtime fully, and reduce overall GPU sharing impact to cloud services. Across a set of ML models, the FIKIT based inference system accelerated high priority tasks by 1.33 to 14.87 times compared to the JCT in GPU sharing mode, and more than half of the cases are accelerated by more than 3.5 times. Alternatively, under preemptive sharing, the low-priority tasks have a comparable to default GPU sharing mode JCT, with a 0.84 to 1 times ratio. We further limit the kernel measurement and runtime fine-grained kernel scheduling overhead to less than 5%.
@@ 4.1. Experimental Setup
      When the B’s low-priority tasks runs continuously in the background, start the A’s high-priority task, the JCT of A’s tasks in the FIKIT mode compared with the GPU sharing mode and the exclusive model.
Hardware and software environments Our experiment used a server with an AMD Ryzen 7 5700X 8-core CPU, 32GB of RAM, and an NVIDIA GeForce RTX 3090 (24GB of device memory). The machine runs Ubuntu 20.04.5 LTS operating system and uses NVIDIA driver version 525.125.06. On the software side, Python3.9, CUDA 11.7, cuDNN 8.8.0, PyTorch 2.0.1, and Torchvision 0.15.2 are used. These configurations constitute the hardware and software environment of our experiments, ensuring the consistency and reproducibility of our experiments.
@@ 4.5.1. JCT Comparison Test (FIKIT VS Share)
We combined services running different models and measure their tasks’ JCT values in FIKIT mode and default sharing mode respectively, and test each combination 1000 times. The task completion time of different models is different. For straight-forward comparison, we take the tasks execution records of the two services overlapping fully. For example, taking two services concurrently running as an example, service A issues keypointrcnn_resnet50_fpn model inference tasks, and service B issues fcn_resnet50 inference tasks. Both services perform 1000 inferences concurrently in FIKIT mode and Share mode respectively. The total time of each service is shown in the table 2. From table 2, it can be concluded that from the beginning of the two services to the 19th second, there are always two concurrent tasks from different services running on the GPU for both experiment settings.
@@
Default GPU Sharing Mode 49.894898 19.72799
FIKIT mode 41.628692 52.227912
@@
Therefore, for this services combination, only the first 19 seconds of JCT data were collected for evaluation. Using the same test method above, 10 combinations of services were tested, and the test results are shown in Fig 16 where ”H” indicates that the model in this mode is a high-priority task, and ”L” indicates that the model in this mode is a low-priority task. In the test cases of 10 services combinations, FIKIT’s JCT is accelerated by 1.33 to 14.87 times compared to the JCT in GPU sharing mode, and more than half of the cases are accelerated by more than 3.5 times, which is a significant performance improvement. If we need both concurrent services serving and guaranteed priority response for some tasks, FIKIT is a good choice.
@@ 4.5.3. High Priority JCT Comparison Test in Preemption Scenario (FIKIT VS Share)
The setting for this test is that service B’s low-priority tasks run continuously, and the service A’s high priority tasks executes intermittently. When B’s low-priority task runs continuously, service A issues a high-priority task every 1 second, and there are in total 100 such tasks. We measure the JCT value of A’s tasks in FIKIT mode and default GPU sharing mode and calculate the average JCT value of 100 tasks. As shown in Fig 19, under most models’ combinations, the JCT of A’s tasks preempting the running GPU tasks under the FIKIT mode is significantly faster than the JCT under the default GPU sharing mode, up to 16.25 times speedup. However, for deeplabv3_resnet50 and resnet101 combination, the FIKIT mode high-priority tasks’ JCT increased compared to the default sharing mode.
@@
As shown in Fig 20, the difference between the JCT values of B’s low-priority tasks in FIKIT mode and default GPU sharing mode is very small, except for the ratio of deeplabv3_resnet50 to resnet101 concurrent tasks, which is only 0.84. The other model combinations’ ratios are very close to each other. Thus, the effect of the FIKIT mode on the operation of service B is almost negligible in this scenario.
@@ 7. Conclusion
Aiming at cloud computing environment, this paper proposed a novel GPU kernel-level preemptive multi-task scheduling architecture, the FIKIT: Filling Inter-kernel Idle Time. Instead of eliminating inter-kernel GPU idle time, we fully utilize these gaps of a time-sensitive GPU task to concurrently execute other sharing GPU tasks, serving multiple tasks having non-symmetric impact on QoS when sharing a GPU device. By a novel mechanism to identify GPU kernels at runtime, the FIKIT find an efficient way to apply offline fine-grained kernel profiling during runtime scheduling with feedback mechanism, separating the profiling impact from GPU multi-task sharing. Across a set of ML models, the FIKIT based inference system accelerated high priority tasks by 1.33 to 14.87 times compared to the job completion time(JCT) in default GPU sharing mode, and more than half of the cases are accelerated by more than 3.5 times. Alternatively, under preemptive sharing, the low-priority tasks have a comparable to default GPU sharing mode JCT, with a 0.84 to 1 times ratio.

### 01 Feb 2024 v5

@@ Abstract.
Highly parallelized workloads like machine learning training, inferences and general HPC tasks are greatly accelerated using GPU devices. In a cloud computing cluster, serving a GPU’s computation power through multi-tasks sharing is highly demanded since there are always more task requests than the number of GPU available. Existing GPU sharing solutions focus on reducing task-level waiting time or task-level switching costs when multiple jobs competing for a single GPU. Non-stopped computation requests come with different priorities, having non-symmetric impact on QoS for sharing a GPU device. Existing work missed the kernel-level optimization opportunity brought by this setting. To address this problem, we present a novel kernel-level scheduling strategy called FIKIT: Filling Inter-kernel Idle Time. FIKIT incorporates task-level priority information, fine-grained kernel identification, and kernel measurement, allowing low priorities task’s execution during high priority task’s inter-kernel idle time. Thereby, filling the GPU’s device runtime fully, and reduce overall GPU sharing impact to cloud services. Across a set of ML models, the FIKIT based inference system accelerated high priority tasks by 1.32 to 16.41 times compared to the JCT in GPU sharing mode, and more than half of the cases are accelerated by more than 3.4 times. Alternatively, under preemptive sharing, the low-priority tasks have a comparable to default GPU sharing mode JCT, with a 0.86 to 1 times ratio. We further limit the kernel measurement and runtime fine-grained kernel scheduling overhead to less than 5%.
@@ 4.1. Experimental Setup
      When service B runs low-priority tasks continuously in the background, A starts high-priority tasks. We evaluate the JCT of A’s tasks in the FIKIT mode compared with the GPU sharing mode and the exclusive model.
    4. (d)
      When service A runs high priority tasks continuously in the background, service B starts the low-priority tasks. We evaluate if B’s tasks JCT is stable and predictable.
Hardware and software environments Our experiment used a server with an AMD Ryzen 7 5700X 8-core CPU, 32GB of RAM, and an NVIDIA GeForce RTX 3090 (24GB of device memory). The machine runs Ubuntu 20.04.5 LTS operating system and uses NVIDIA driver version 525.147.05 (525.125.06 is also tested with all the experiments while not included in this paper). On the software side, Python3.9, CUDA 11.7 , cuDNN 8.8.0, PyTorch 2.0.1, and Torchvision 0.15.2 are used. These configurations constitute the hardware and software environment of our experiments, ensuring the consistency and reproducibility of our experiments.
@@ 4.5.1. JCT Comparison Test (FIKIT VS Share)
We combined services running different models and measure their tasks’ JCT values in FIKIT mode and default sharing mode respectively, and test each combination 1000 times. The task completion time of different models is different. For straight-forward comparison, we take the tasks execution records of the two services overlapping fully. For example, taking two services concurrently running as an example, service A issues keypointrcnn_resnet50_fpn model inference tasks, and service B issues fcn_resnet50 inference tasks. Both services perform 1000 inferences concurrently in FIKIT mode and Share mode respectively. The total time of each service is shown in the table 2. From table 2, it can be concluded that from the beginning of the two services to the 16th second, there are always two concurrent tasks from different services running on the GPU for both experiment settings.
@@
Default GPU Sharing Mode 38.15803 16.02363
FIKIT mode 33.12627 39.09784
@@
Therefore, for this services combination, only the first 16 seconds of JCT data were collected for evaluation. Using the same test method above, 10 combinations of services were tested, and the test results are shown in Fig 16 where ”H” indicates that the model in this mode is a high-priority task, and ”L” indicates that the model in this mode is a low-priority task. In the test cases of 10 services combinations, FIKIT’s JCT is accelerated by 1.32 to 16.41 times compared to the JCT in GPU sharing mode, and more than half of the cases are accelerated by more than 3.4 times, which is a significant performance improvement. If we need both concurrent services serving and guaranteed priority response for some tasks, FIKIT is a good choice.
@@ 4.5.3. High Priority JCT Comparison Test in Preemption Scenario (FIKIT VS Share)
The setting for this test is that service B’s low-priority tasks run continuously, and the service A’s high priority tasks executes intermittently. When B’s low-priority task runs continuously, service A issues a high-priority task every 1 second, and there are in total 100 such tasks. We measure the JCT value of A’s tasks in FIKIT mode and default GPU sharing mode and calculate the average JCT value of 100 tasks. As shown in Fig 19, under most models’ combinations, the JCT of A’s tasks preempting the running GPU tasks under the FIKIT mode is significantly faster than the JCT under the default GPU sharing mode, up to 15.77 times speedup. However, for deeplabv3_resnet50 and resnet101 combination, the FIKIT mode high-priority tasks’ JCT increased compared to the default sharing mode.
@@
As shown in Fig 20, the difference between the JCT values of B’s low-priority tasks in FIKIT mode and default GPU sharing mode is very small, except for the ratio of deeplabv3_resnet50 to resnet101 concurrent tasks, which is only 0.86. The other model combinations’ ratios are very close to each other. Thus, the effect of the FIKIT mode on the operation of service B is almost negligible in this scenario.
@@
4.5.4. Low Priority JCT Stability Test in FIKIT Sharing Scenario
The setting for this test is that service A runs high-priority tasks continuously, and service B executes low-priority tasks intermittently. In detail, service B issues a low-priority task every 1 second, and there are in total 100 such tasks. We measure the JCT value of B’s tasks in FIKIT mode and calculate the coefficient of variation to evaluate low-priority tasks JCT stability spanning 100 seconds. As shown in Fig 21, under every models’ combinations, the JCT of low-priority tasks scavenging the inter-kenel idling time of high-priority GPU tasks under the FIKIT sharing mode is stable and predictable. We also show that coefficient of variations are within 0.095 to 0.164 for 10 combinations of models spanning 100 seconds in Table 3. These timelines’ CV being way smaller than 1 indicates a low variability and high predictability of low-priority tasks’ job completion time when there are continuously runing high-priority tasks using FIKIT sharing mode.
The stability and predictable JCT for low-priority tasks is a desired guarantee in cloud computing environement and we show that FIKIT have achieved this guarantee under multi-task GPU sharing scenario.
Figure 21. Low Priority JCT time lines under FIKIT sharing
Timeline $\sigma$ (ms) $\mu$ (ms) CV = $\frac{\sigma}{\mu}$
A 10.047 61.391 0.163657
B 18.428 176.578 0.104361
C 11.948 104.195 0.114673
D 8.692 55.408 0.156871
E 15.962 127.269 0.125423
F 0.670 7.017 0.095457
G 10.524 77.409 0.135951
H 21.547 136.410 0.157956
I 12.393 114.532 0.108204
J 4.616 41.719 0.110651
Table 3. Standard Deviation, Mean, and Coefficient of Variance for each Timeline
@@ 7. Conclusion
Aiming at cloud computing environment, this paper proposed a novel GPU kernel-level preemptive multi-task scheduling architecture, the FIKIT: Filling Inter-kernel Idle Time. Instead of eliminating inter-kernel GPU idle time, we fully utilize these gaps of a time-sensitive GPU task to concurrently execute other sharing GPU tasks, serving multiple tasks having non-symmetric impact on QoS when sharing a GPU device. By a novel mechanism to identify GPU kernels at runtime, the FIKIT find an efficient way to apply offline fine-grained kernel profiling during runtime scheduling with feedback mechanism, separating the profiling impact from GPU multi-task sharing. Across a set of ML models, the FIKIT based inference system accelerated high priority tasks by 1.32 to 16.41 times compared to the job completion time(JCT) in default GPU sharing mode, and more than half of the cases are accelerated by more than 3.4 times. Alternatively, under preemptive sharing, the low-priority tasks have a comparable to default GPU sharing mode JCT, with a 0.86 to 1 times ratio.